OpenStudy (anonymous):

8(squared) x 90(square) multiply and simplify by factoring

OpenStudy (saifoo.khan):

\[\Large 8^2 \times 90^2 \]

OpenStudy (saifoo.khan):

That?

OpenStudy (anonymous):

\[\sqrt{8}\times \sqrt{90}\]

OpenStudy (anonymous):

not squared the square root symbol sorry

OpenStudy (saifoo.khan):

\[\sqrt{ 8 * 90}\]Solve.

OpenStudy (anonymous):

I thing it 720? \[\sqrt{720}\]

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

But if you simplify it, it becomes\[12\sqrt{5}\]